Calculate the molecular weight of an ethylene molecule with a degree of polymerization of 500.
Answer #1

given that

degree of polymerization Xn =500

the molecular formula of ethylene = c2H4

the molecular weight of ethylene =M=2[12.01 g/mol]+4[1.008 g/mol] =28.052 g/mol

the degree of polymerization Xn = Mn/m

therefore the average molecular weight Mn=Xn*m

=500*28.052=14026 g/mol